# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

### v2

@@ Abstract
$\qquad\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})\big{)}}\Bigg{)},$ (S40a)
@@
$\qquad-\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})\big{)}}\Bigg{)},$ (S40b)
@@
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{2}+\text{s}k_{y}^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{2}(\bm{k})+\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{2}(\bm{k})))}{(\mu+E_{2}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k}))},$
@@
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{8}+\text{s}k_{y}^{8})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{8}(\bm{k})+\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{8}(\bm{k})))}{(\mu+E_{8}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})+e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})}\Bigg{)},$ (S41a)
@@
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{2}+\text{s}k_{y}^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{4}(\bm{k})-\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{4}(\bm{k})))}{(\mu+E_{4}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k}))},$
@@
$\frac{2A\mu(i\text{s}(k_{x})+\text{s}(k_{y}))}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu(\text{s}k_{x}^{2}+\text{s}k_{y}^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{6}(\bm{k})-\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{6}(\bm{k})))}{(\mu+E_{6}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})-e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})}\Bigg{)}.$ (S41b)
At least one component of each of the $\Psi_{n,\bm{k}}$ in (S40,S41) is nonzero at every $\bm{k}\in\text{BZ}$, thus $\forall\bm{k}\in\text{BZ}:\innerproduct{\Psi_{n,\bm{k}}}{\Psi_{n,\bm{k}}}\neq 0$. Indeed for every $\bm{k}\in\text{BZ}$ each of the vectors |$\ket{\Psi_{1,\bm{k}}},\ldots,\ket{\Psi_{8,\bm{k}}}$ defined by Eq. (S40,S41) are energy eigenvectors of Bloch’s form (they satisfy |$\ket{\Psi_{n,\bm{k}+\bm{G}}}=\ket{\Psi_{n,\bm{k}}}$) and are smooth over BZ. These vectors can be normalized to obtain
|$\ket{\psi_{n,\bm{k}}}\equiv\frac{\ket{\Psi_{n,\bm{k}}}}{\innerproduct{\Psi_{n,\bm{k}}}{\Psi_{n,\bm{k}}}}.$ (S42)
@@
We identify the components of the non-Abelian Berry connection induced by the gauge $\mathfrak{u}$ by
$\xi^{a}_{nm}(\bm{k})\equiv\big{(}\xi^{\mathfrak{u}}(\bm{k})\big{)}^{a}_{nm}=i\innerproduct{u_{n,\bm{k}}}{\partial_{a}u_{m,\bm{k}}}.$ (S43)
In principle, the components (S43) can be obtained by explicitly taking $\bm{k}$-derivatives of Eqs. (S40,S41) and employing some approximation for the $\bm{k}$-derivatives of the |$\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}$ (for example, using the assumption of atomic-like WFs or of highly-localized WFs as described above). However, we will perform our calculations numerically. Some useful (numerically obtained) relations that arise in this gauge choice are
@@
$\xi^{a}_{35}=-\xi^{a}_{64},\,\xi^{a}_{36}=\xi^{a}_{54},\,\xi^{a}_{37}=-\xi^{a}_{84},\,\xi^{a}_{38}=\xi^{a}_{74}.$ (S44)
The relations (S44) arise in both the approximation of atomic-like WFs ($\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$) and the approximation of highly-localized WFs (S29) for $\bm{\delta}_{2}=\bm{\delta}_{3}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=\bm{\delta}_{1}=-\frac{d}{2}\bm{z}$. These relations are analogous to those presented in Eq. (42) of Ref. [15] for the non-magnetic case. Indeed each of the relations in Eq. (S44) have a natural analog in the non-magnetic case. However, in the non-magnetic case there are relations that have no analog here; in the Hamiltonian gauge employed in the non-magnetic case [15], $\xi^{a}_{vv^{\prime}}\propto\xi^{a}_{cc^{\prime}}$ for various combinations of $v$, $v^{\prime}$, $c$, and $c^{\prime}$ (in particular, $\xi^{a}_{11}=\xi^{a}_{33}$, $\xi^{a}_{22}=\xi^{a}_{44}$ and $\xi^{a}_{12}=-\xi^{a}_{34}$, where bands 1 and 2 are filled and bands 3 and 4 are empty). In the AFM case we do not have a relation between any $\xi^{a}_{vv}$ and $\xi^{a}_{cc}$, nor do we have a relation between $\xi^{a}_{12}$ and $\xi^{a}_{78}$ or between $\xi^{a}_{34}$ and $\xi^{a}_{56}$. As we later demonstrate, the lack of such relations will result in a qualitative distinction between the partitioning of the magnetoelectric response into atomic-like and itinerant contributions in the non-magnetic and AFM cases.
In the AFM case there also emerges a property of certain $\xi^{a}_{nm}$ that is not present in the non-magnetic case. It turns out that certain band components of the AFM $\xi^{a}$ are numerically non-negligible only when $a=z$. In particular,
$\xi^{a}_{13},\,\xi^{a}_{14},\,\xi^{a}_{15},\,\xi^{a}_{16},\,\xi^{a}_{23},\,\xi^{a}_{24},\,\xi^{a}_{25},\,\xi^{a}_{26},\,\xi^{a}_{73},\,\xi^{a}_{74},\,\xi^{a}_{75},\,\xi^{a}_{76},\,\xi^{a}_{83},\,\xi^{a}_{84},\,\xi^{a}_{85},\,\xi^{a}_{86}$ $\neq 0\text{ only if }a=3.$ (S45)
S2.2 Computation of $\alpha_{\text{CS}}$
@@
$\alpha^{\mathfrak{u}}_{\text{CS}}=-\frac{e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\left(\sum_{vv^{\prime}}\xi^{a}_{vv^{\prime}}(\bm{k})\partial_{b}\xi^{d}_{v^{\prime}v}(\bm{k})-\frac{2i}{3}\sum_{vv^{\prime}v_{1}}\xi^{a}_{vv^{\prime}}(\bm{k})\xi^{b}_{v^{\prime}v_{1}}(\bm{k})\xi^{d}_{v_{1}v}(\bm{k})\right),$ (S46)
where $v,v^{\prime},v_{1}\in\{1,2,3,4\}$ in this case. Again, the gauge $\mathfrak{u}$ is defined by Eq. (S40,S41). We evaluate this integral numerically and interpolate to convergence. Some results are listed in Table 1. To correspond with MBX materials we focus on the parameter regime $B/\Delta_{\text{S}}<0$, and arbitrarily choose $\Delta_{\text{S}}=190$ meV and $B=-25$ meV. The values of $\alpha^{\mathfrak{u}}_{\text{CS}}$ listed in Table 1 are independent of $A\in\mathbb{R}_{+}$, which is a parameter that is related to the single layer lattice constant that was introduced during regularization of the effective low-energy coupled Dirac cone model. Notably, when the values of the crystal parameters are varied such that the pairs of degenerate bands remain isolated from one another, the value to which $\alpha^{\mathfrak{u}}_{\text{CS}}$ evaluates is unchanged. The findings listed in Table 1 are consistent with Fig. 2 (a) of the main text. Due to the large number of parameters in our model (Eq. (S38)) we do not extract the entire topological phase diagram. Via explicit calculation we have found that the value to which $\alpha^{\mathfrak{u}}_{\text{CS}}$ evaluates is unchanged regardless of whether we adopt either of the above described approximations of atomic-like WFs ($\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$) or of highly-localized WFs (S29) with $\bm{\delta}_{2}=\bm{\delta}_{3}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=\bm{\delta}_{1}=-\frac{d}{2}\bm{z}$.
@@
Table 1: Values of $\alpha^{\mathfrak{u}}_{\text{CS}}$ obtained from numerical integration of Eq. (S46) for $\Delta_{\text{S}}=190$ meV and $B=-25$ meV.
S2.3 Partitioning of $\alpha_{\text{CS}}$ into atomic-like and itinerant contributions
S2.3.1 Considerations for a general smooth global Hamiltonian gauge
We begin with a discussion of the linear response of orbital magnetization in a band insulator that exhibits TRS. We then restrict focus to our model, Eq. (S38). In general, the electronic orbital magnetization $\bm{M}$ can be partitioned into a sum of atomic-like $\overline{\bm{M}}$ and itinerant $\widetilde{\bm{M}}$ contributions,
$\bm{M}=\overline{\bm{M}}+\widetilde{\bm{M}}.$ (S47)

### v4

@@ B.1 Tight-binding model
$\qquad\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{7}(\bm{k})\big{)}}\Bigg{)},$ (49a)
@@
$\qquad-\frac{2e^{2idk_{z}}A\sqrt{\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2}}(-i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu-(1+e^{2idk_{z}})E_{5}(\bm{k})\big{)}}\Bigg{)},$ (49b)
@@
$\frac{2A\mu(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu((\text{s}k_{x})^{2}+(\text{s}k_{y})^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{2}(\bm{k})+\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{2}(\bm{k})))}{(\mu+E_{2}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{2}(\bm{k}))},$
@@
$\frac{2A\mu(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu((\text{s}k_{x})^{2}+(\text{s}k_{y})^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{8}(\bm{k})+\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{8}(\bm{k})))}{(\mu+E_{8}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})+e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{8}(\bm{k})}\Bigg{)},$ (50a)
@@
$\frac{2A\mu(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu((\text{s}k_{x})^{2}+(\text{s}k_{y})^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{4}(\bm{k})-\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{4}(\bm{k})))}{(\mu+E_{4}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{4}(\bm{k}))},$
@@
$\frac{2A\mu(i\text{s}k_{x}+\text{s}k_{y})}{\Delta_{\text{D}}\big{(}(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})\big{)}},$
$\frac{((-1+e^{2idk_{z}})\Delta_{\text{D}}^{2}\mu-2A^{2}\mu((\text{s}k_{x})^{2}+(\text{s}k_{y})^{2})-2\mu\Delta_{\text{S}}(k_{x},k_{y})^{2}+e^{idk_{z}}\Delta_{\text{D}}^{2}\text{c}(dk_{z})2E_{6}(\bm{k})-\sqrt{(\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(dk_{z})^{2})\Delta_{\text{S}}(k_{x},k_{y})^{2}}(-2\mu+2E_{6}(\bm{k})))}{(\mu+E_{6}(\bm{k}))\Delta_{\text{D}}((-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k}))},$
$-\frac{(1+e^{2idk_{z}})A(-i\text{s}k_{x}-\text{s}k_{y})}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})},\,1,\,0,\,\frac{(1+e^{2idk_{z}})\Delta_{\text{s}}(k_{x},k_{y})-e^{2idk_{z}}\sqrt{2(\Delta_{\text{D}}^{2}+2\mu^{2}+\Delta_{\text{D}}^{2}\text{c}(2dk_{z}))}}{(-1+e^{2idk_{z}})\mu+(1+e^{2idk_{z}})E_{6}(\bm{k})}\Bigg{)}.$ (50b)
At least one component of each of the $\Psi_{n,\bm{k}}$ in Eqs. (49) and (50) is nonzero at every $\bm{k}\in\text{BZ}$, thus $\forall\bm{k}\in\text{BZ}:\innerproduct{\Psi_{n,\bm{k}}}{\Psi_{n,\bm{k}}}\neq 0$. Indeed for every $\bm{k}\in\text{BZ}$ each of the vectors |$\ket{\Psi_{1,\bm{k}}},\ldots,\ket{\Psi_{8,\bm{k}}}$ defined by Eqs. (49) and (50) are energy eigenvectors of Bloch’s form (they satisfy |$\ket{\Psi_{n,\bm{k}+\bm{G}}}=\ket{\Psi_{n,\bm{k}}}$) and are smooth over BZ. These vectors can be normalized to obtain
|$\ket{\psi_{n,\bm{k}}}\equiv\frac{\ket{\Psi_{n,\bm{k}}}}{\innerproduct{\Psi_{n,\bm{k}}}{\Psi_{n,\bm{k}}}}.$ (51)
@@
We identify the components of the nonAbelian Berry connection induced by the gauge $\mathfrak{u}$ by
$\xi^{a}_{nm}(\bm{k})\equiv\big{(}\xi^{\mathfrak{u}}(\bm{k})\big{)}^{a}_{nm}=i\innerproduct{u_{n,\bm{k}}}{\partial_{a}u_{m,\bm{k}}}.$ (52)
In principle, the components (52) can be obtained by explicitly taking $\bm{k}$-derivatives of Eqs. (49) and (50) and employing some approximation for the $\bm{k}$-derivatives of the |$\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}$ (for example, using the assumption of atomic-like WFs or of highly localized WFs as described above). However, we will perform our calculations numerically. Some useful (numerically obtained) relations that arise in this gauge choice are
@@
$\xi^{a}_{35}=-\xi^{a}_{64},\,\xi^{a}_{36}=\xi^{a}_{54},\,\xi^{a}_{37}=-\xi^{a}_{84},\,\xi^{a}_{38}=\xi^{a}_{74}.$ (53)
The relations (53) arise in both the approximation of atomic-like WFs ($\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$) and the approximation of highly localized WFs (38) for $\bm{\delta}_{2}=\bm{\delta}_{3}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=\bm{\delta}_{1}=-\frac{d}{2}\bm{z}$. These relations are analogous to those presented in Eq. (42) of Ref. [15] for the nonmagnetic case. Indeed each of the relations in Eq. (53) have a natural analog in the nonmagnetic case. However, in the nonmagnetic case there are relations that have no analog here; in the Hamiltonian gauge employed in the nonmagnetic case [15], $\xi^{a}_{vv^{\prime}}\propto\xi^{a}_{cc^{\prime}}$ for various combinations of $v$, $v^{\prime}$, $c$, and $c^{\prime}$ (in particular, $\xi^{a}_{11}=\xi^{a}_{33}$, $\xi^{a}_{22}=\xi^{a}_{44}$ and $\xi^{a}_{12}=-\xi^{a}_{34}$, where bands 1 and 2 are filled and bands 3 and 4 are empty). In the AFM case we do not have a relation between any $\xi^{a}_{vv}$ and $\xi^{a}_{cc}$, nor do we have a relation between $\xi^{a}_{12}$ and $\xi^{a}_{78}$ or between $\xi^{a}_{34}$ and $\xi^{a}_{56}$. As we later demonstrate, the lack of such relations will result in a qualitative distinction between the partitioning of the magnetoelectric response into atomic-like and itinerant contributions in the nonmagnetic and AFM cases.
In the AFM case there also emerges a property of certain $\xi^{a}_{nm}$ that is not present in the nonmagnetic case. It turns out that certain band components of the AFM $\xi^{a}$ are numerically nonnegligible only when $a=z$. In particular,
$\xi^{a}_{13},\,\xi^{a}_{14},\,\xi^{a}_{15},\,\xi^{a}_{16},\,\xi^{a}_{23},\,\xi^{a}_{24},\,\xi^{a}_{25},\,\xi^{a}_{26},\,\xi^{a}_{73},\,\xi^{a}_{74},\,\xi^{a}_{75},\,\xi^{a}_{76},\,\xi^{a}_{83},\,\xi^{a}_{84},\,\xi^{a}_{85},\,\xi^{a}_{86}$ $\neq 0\text{ only if }a=3.$ (54)
B.2 Computation of $\alpha_{\text{CS}}$
@@
$\alpha^{\mathfrak{u}}_{\text{CS}}=-\frac{e^{2}}{2\hbar c}\epsilon^{abd}\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\left(\sum_{vv^{\prime}}\xi^{a}_{vv^{\prime}}(\bm{k})\partial_{b}\xi^{d}_{v^{\prime}v}(\bm{k})-\frac{2i}{3}\sum_{vv^{\prime}v_{1}}\xi^{a}_{vv^{\prime}}(\bm{k})\xi^{b}_{v^{\prime}v_{1}}(\bm{k})\xi^{d}_{v_{1}v}(\bm{k})\right),$ (55)
where $v,v^{\prime},v_{1}\in\{1,2,3,4\}$ in this case. Again, the gauge $\mathfrak{u}$ is defined by Eqs. (49) and (50). We evaluate this integral numerically and interpolate to convergence. Some results are listed in Table 1. To correspond with MBX materials we focus on the parameter regime $B/\Delta_{\text{S}}<0$, and arbitrarily choose $\Delta_{\text{S}}=190$ meV and $B=-25$ meV. The values of $\alpha^{\mathfrak{u}}_{\text{CS}}$ listed in Table 1 are independent of $A\in\mathbb{R}_{+}$, which is a parameter that is related to the single layer lattice constant that was introduced during regularization of the effective low-energy coupled Dirac cone model. Notably, when the values of the crystal parameters are varied such that the pairs of degenerate bands remain isolated from one another, the value to which $\alpha^{\mathfrak{u}}_{\text{CS}}$ evaluates is unchanged. The findings listed in Table 1 are consistent with Fig. 2 (a) of the main text. Due to the large number of parameters in our model (Eq. (47)) we do not extract the entire topological phase diagram. Via explicit calculation we have found that the value to which $\alpha^{\mathfrak{u}}_{\text{CS}}$ evaluates is unchanged regardless of whether we adopt either of the above described approximations of atomic-like WFs ($\partial_{a}\ket{\bar{u}_{(\alpha,\sigma),\bm{k}}}\equiv 0$) or of highly localized WFs (38) with $\bm{\delta}_{2}=\bm{\delta}_{3}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=\bm{\delta}_{1}=-\frac{d}{2}\bm{z}$.
@@
Table 1: Values of $\alpha^{\mathfrak{u}}_{\text{CS}}$ obtained from numerical integration of Eq. (55) for $\Delta_{\text{S}}=190$ meV and $B=-25$ meV.
B.3 Partitioning of $\alpha_{\text{CS}}$ into atomic-like and itinerant contributions
B.3.1 Considerations for a general smooth global Hamiltonian gauge
We begin with a discussion of the linear response of orbital magnetization in a band insulator that exhibits TRS. We then restrict focus to our model, Eq. (47). In general, the electronic orbital magnetization $\bm{M}$ can be partitioned into a sum of atomic-like $\overline{\bm{M}}$ and itinerant $\widetilde{\bm{M}}$ contributions,
$\bm{M}=\overline{\bm{M}}+\widetilde{\bm{M}}.$ (56)